OpenStudy (anonymous):

Convert 900 milligrams into grams.

OpenStudy (pounce0129):

0.9 grams

hero (hero):

900 mg = x g Divide both sides by g to get 900 m = x Change m to 1/1000 900/1000 = x 9/10 = x 0.9 = x

OpenStudy (anonymous):

1000 milligrams = 1 gram 1 milligram = 1/1000 grams 900 milligrams = (1/1000) *900 grams 900 milligrams = 0.9 grams

hero (hero):

My strategy doesn't require memorizing how many milligrams in a gram

OpenStudy (anonymous):

@kiki41554 got it..??